# Resonant W and Z Boson Production in FSRQ Jets: Implications for Diffuse Neutrino Fluxes

### v2

@@ 3.3 Diffuse neutrino flux from electron–positron collisions in FSRQ jets
and parameters $z_{c}^{*}=1.47$, $\kappa=0.21$, $p_{1}=7.35$, $p_{2}=-6.51$. The event rate $\Gamma_{e^{-}e^{+}}^{j}(L)$ of electron–positron collisions producing bosons (i.e., $j=W^{\pm},Z$) is assumed to scale with the blazar luminosity $L$ as
$\Gamma_{e^{-}e^{+}}^{j}(L)\sim\Gamma_{e^{-}e^{+},0}^{j}\frac{L}{L_{0}},$ (3.13)
where $L_{0}$ and $\Gamma_{e^{-}e^{+},0}^{j}$ are the luminosity and corresponding event rate for the reference source 3C 279. This scaling is justified by the expectation that the density of high-energy electrons and positrons, and hence the interaction rate, should increase with the total nonthermal power output of the jet, which is commonly traced by the bolometric luminosity. For a source at redshift $z$, the observed flux $F_{e^{-}e^{+}}^{j,\rm obs}(L,z)$ is given by
@@
where $dV_{c}/dzd\Omega$ is the comoving volume element per unit redshift and solid angle, obtained from the cosmological model [50]. The redshift and source luminosity ranges are assumed to be $z=[0.01,5]$ and $L=[10^{44},10^{50}]\,\mathrm{erg\,s^{-1}}$, respectively, covering the typical luminosity and redshift distribution of FSRQs. To investigate the flux contribution from different cosmic epochs, the integrands of $F_{e^{-}e^{+}}^{Z,\rm diffuse}$ are shown in figure 7-(a) as functions of luminosity $L$ at various redshifts $z$. As expected from the form of the luminosity function, sources at $z>2$ provide substantial contributions when their luminosities exceed $10^{48}\,\mathrm{erg\,s^{-1}}$. Integrating over luminosity yields the differential flux contribution, ${dF_{e^{-}e^{+}}^{Z,\rm diffuse}}/{dz}$, which is estimated as a function of redshift $z$ (figure 7-(b)), exhibiting a clear peak around $z\sim 1$. This peak reflects the cosmic evolution of the FSRQ population, where the comoving number density of luminous sources reaches its maximum. At lower redshifts the volume density decreases, while at higher redshifts the increasing luminosity distance and declining source density reduce the overall contribution. Therefore, the differential flux peak around $z\sim 1$ represents a physically meaningful imprint of the evolutionary history of FSRQs.
@@
$F_{e^{-}e^{+}}^{W^{\pm},\rm diffuse}$ $\sim 2.7\times 10^{-26}\ \mathrm{cm^{-2}\,s^{-1}\,sr^{-1}},$ (3.16)
$F_{e^{-}e^{+}}^{Z,\rm diffuse}$ $\sim 1.1\times 10^{-15}\ \mathrm{cm^{-2}\,s^{-1}\,sr^{-1}}.$ (3.17)
The estimated fluxes represent the cumulative contribution of electron–positron annihilation into $W^{\pm}$ and $Z$ bosons across the cosmological population of FSRQs. Although the obtained values remain well below the diffuse astrophysical neutrino flux measured by current high-energy neutrino observatories ($\sim 10^{-12}\ \mathrm{cm^{-2}\,s^{-1}\,sr^{-1}}$ at 0.1–1 TeV) such as IceCube [e.g., 45] or KM3NeT [e.g., 46], they provide a theoretical benchmark for the role of leading (resonant) Standard Model processes in extreme astrophysical environments. This result highlights the connection between particle physics and astrophysics, demonstrating that even rare high-energy interactions leave a subtle but well-defined imprint on the cosmic radiation background. While the predicted flux is currently undetectable, it establishes an upper limit for diffuse electroweak signatures. In particular, the resonant $Z$ boson production channel yields a flux that is about $\sim 10^{-3}$ of the total diffuse astrophysical neutrino flux, representing a small yet theoretically meaningful fraction that serves as a probe of high-energy electroweak processes in FSRQ jets. This interpretation remains valid when considering the astrophysical uncertainties involved in modeling the electron energy distribution in the jet blob, including acceleration and energy losses. Since the reaction rate may vary by about an order of magnitude as shown in figure 5, such uncertainties are less significant when compared to the large gap between the diffuse neutrino flux expected from resonant $Z$ boson production and the sensitivity of currently available neutrino detectors. In addition, the predicted fluxes can, in principle, be modified (for example, through enhanced cross sections) by incorporating new mechanisms of particle production in lepton collisions beyond the Standard Model. The presented framework readily allows one to take into account such modifications.
@@ 4 Summary and discussion
We employ a one-zone leptonic model to describe the jet dynamics and electron distribution in the FSRQ 3C 279 during its flaring state, using the Fokker–Planck equation to model electron acceleration via shock processes and turbulence. This model allows us to estimate the reaction rates for the production of $W^{\pm}$ and $Z$ bosons from electron–positron collisions. The results show that while these processes are theoretically possible, the flux from such interactions is currently well below the sensitivity limits of existing neutrino observatories such as IceCube, KM3NeT, or Baikal-GVD. Additionally, the estimated diffuse flux from the cosmological population of FSRQs suggests a significant contribution from $Z$ boson production, which could account for approximately $\sim 10^{-3}$ of the total diffuse astrophysical neutrino flux.
Despite the low detectability of these signals, the study highlights the importance of investigating electroweak interactions in extreme astrophysical environments. Since the modeling is based on the flaring state of 3C 279, the estimated interaction rates should be regarded as optimistic upper limits; quiescent states would likely yield even smaller contributions. Although the predicted fluxes remain beyond the reach of current observatories, future facilities such as IceCube-Gen2 or KM3NeT may begin to test such scenarios, especially in synergy with multi-messenger observations of blazar flares. Our simplified one-zone framework provides a first benchmark, while more elaborate models including multi-zone dynamics, hadronic processes, or time-dependent variability will further improve the estimates.
@@ Acknowledgments
The authors thank S. V. Troitsky and D. S. Gorbunov for useful discussions.

### v3

@@ 3.3 Diffuse neutrino flux from electron–positron collisions in FSRQ jets
and parameters $z_{c}^{*}=1.47$, $\kappa=0.21$, $p_{1}=7.35$, $p_{2}=-6.51$. The event rate $\Gamma_{e^{-}e^{+}}^{j}(L)$ of electron–positron collisions producing electroweak bosons ($j=W^{\pm},Z$) is assumed to scale with the blazar luminosity $L$ according to
$\Gamma_{e^{-}e^{+}}^{j}(L)=\Gamma_{e^{-}e^{+},0}^{j}\left(\frac{L}{L_{0}}\right)^{2},$ (3.13)
where $L_{0}$ and $\Gamma_{e^{-}e^{+},0}^{j}$ denote the luminosity and corresponding event rate of the reference source 3C 279. In the optically thin regime the interaction rate scales as the product of the electron and positron number densities $n_{e^{-}}n_{e^{+}}$. If the densities of both energetic electrons and positrons increase with the jet power traced by $L$ a quadratic dependence $\Gamma\propto L^{2}$ naturally follows. This assumes that both electron and positron densities scale with the jet power in a similar manner, as expected in pair rich FSRQ jets where the lepton content is constrained by SED modeling. This scaling therefore represents a physically motivated extrapolation to the blazar population reflecting the expectation that the density of high energy leptons and hence the collision rate grows with the total nonthermal power output of the jet which is commonly traced by the bolometric luminosity. For a source at redshift $z$, the observed flux $F_{e^{-}e^{+}}^{j,\rm obs}(L,z)$ is given by
@@
where $dV_{c}/dzd\Omega$ is the comoving volume element per unit redshift and solid angle, obtained from the cosmological model [51]. The redshift and source luminosity ranges are assumed to be $z=[0.01,5]$ and $L=[10^{44},10^{50}]\,\mathrm{erg\,s^{-1}}$, respectively, covering the typical luminosity and redshift distribution of FSRQs. To investigate the flux contribution from different cosmic epochs, the integrands of $F_{e^{-}e^{+}}^{Z,\rm diffuse}$ are shown in figure 7-(a) as functions of luminosity $L$ at various redshifts $z$. As expected from the form of the luminosity function, sources at $z>2$ provide substantial contributions when their luminosities exceed $10^{48}\,\mathrm{erg\,s^{-1}}$. Integrating over luminosity yields the differential flux contribution, ${dF_{e^{-}e^{+}}^{Z,\rm diffuse}}/{dz}$, which is estimated as a function of redshift $z$ (figure 7-(b)), exhibiting a clear peak around $z\sim 1.5$. This peak reflects the cosmic evolution of the FSRQ population, where the comoving number density of luminous sources reaches its maximum. At lower redshifts the volume density decreases, while at higher redshifts the increasing luminosity distance and declining source density reduce the overall contribution. Therefore, the differential flux peak around $z\sim 1.5$ represents a physically meaningful imprint of the evolutionary history of FSRQs.
@@
$F_{e^{-}e^{+}}^{W^{\pm},\rm diffuse}$ $\sim 3.7\times 10^{-36}\ \mathrm{cm^{-2}\,s^{-1}\,sr^{-1}},$ (3.16)
$F_{e^{-}e^{+}}^{Z,\rm diffuse}$ $\sim 1.5\times 10^{-25}\ \mathrm{cm^{-2}\,s^{-1}\,sr^{-1}}.$ (3.17)
The estimated fluxes represent the cumulative contribution of electron–positron annihilation into $W^{\pm}$ and $Z$ bosons across the cosmological population of FSRQs. Although the obtained values remain well below the diffuse astrophysical neutrino flux measured by current high-energy neutrino observatories ($\sim 10^{-12}\ \mathrm{cm^{-2}\,s^{-1}\,sr^{-1}}$ at 0.1–1 TeV) such as IceCube [e.g., 46] or KM3NeT [e.g., 47], they provide a theoretical benchmark for the role of leading (resonant) Standard Model processes in extreme astrophysical environments. This result highlights the connection between particle physics and astrophysics, demonstrating that even extremely rare high-energy interactions leave a subtle imprint on the cosmic radiation background. While the predicted flux is far below current detection capabilities, it provides a theoretical reference for diffuse electroweak signatures. In particular, when considering interactions confined within the jet blob, the resonant $Z$ boson production channel yields a flux that is many orders of magnitude smaller than the total diffuse astrophysical neutrino flux. Despite its extreme faintness, this fraction remains of theoretical interest as a probe of high-energy electroweak processes in FSRQ jets. This interpretation continues to hold under the astrophysical uncertainties associated with modeling the electron energy distribution within the blob, including acceleration and energy losses. Since the reaction rate may vary by about an order of magnitude as shown in figure 5, such uncertainties are less significant when compared to the large gap between the diffuse neutrino flux expected from resonant $Z$ boson production and the sensitivity of currently available neutrino detectors. In addition, the predicted fluxes can, in principle, be modified (for example, through enhanced cross sections) by incorporating new mechanisms of particle production in lepton collisions beyond the Standard Model. The presented framework readily allows one to take into account such modifications.
@@ 4 Summary and discussion
We employ a one-zone leptonic model to describe the jet dynamics and electron distribution in the FSRQ 3C 279 during its flaring state, using the Fokker–Planck equation to model electron acceleration via shock processes and turbulence. This model allows us to estimate the reaction rates for the production of $W^{\pm}$ and $Z$ bosons from electron–positron collisions confined within the jet blob. The results show that while these processes are theoretically possible, the flux from such interactions is many orders of magnitude below the sensitivity limits of existing neutrino observatories such as IceCube, KM3NeT, or Baikal-GVD. Consequently, the contribution from $Z$ boson production to the diffuse astrophysical neutrino flux is extremely small, far below previous estimates.
Despite the low detectability of these signals, the study highlights the importance of investigating electroweak interactions in extreme astrophysical environments. Since the modeling is based on the flaring state of 3C 279, the estimated interaction rates should be regarded as optimistic upper limits; quiescent states would likely yield even smaller contributions. Although the predicted fluxes remain beyond the reach of current observatories, future facilities such as IceCube-Gen2 or KM3NeT may begin to explore such scenarios, especially in synergy with multi-messenger observations of blazar flares. Our simplified one-zone framework provides a first theoretical benchmark, while more elaborate models including multi-zone dynamics, hadronic processes, or time-dependent variability will further refine the estimates.
@@ Acknowledgments
The authors thank S. V. Troitsky and D. S. Gorbunov for useful discussions. We also thank the referees for their careful reading of the manuscript and for their constructive comments and suggestions, which have significantly improved the quality and clarity of this work.